sir, a bubble in glass slab of refractive index 1.5 when viewed from one side appears at 5 cm and 2cm from other side then thickness of glass slab is.

sir,
a bubble in glass slab of refractive index 1.5 when viewed from one side appears at 5 cm and 2cm from other side then thickness of glass slab is.

Grade:

2 Answers

Sayandip nath
6 years ago
Let is take d1= 5cm and d2=2cm and x=refractive index=1.5Now. Actual width is the sum of real depth from the 2 sides= d1*x + d2*x= 2*1.5 + 5*1.5= 10.5cm
Jay
6 years ago
Acco. To given information. [ #] i use this letter for refractive index we know refractive index = real depth/ apparent depth case 1. x/5=1.5 and case 2 x/2= 1.5 now add both case we get x = 10.5 as answer